OpenStudy (anonymous):

solve for x. show all steps leading to the solution 3(x-1)=2x-3+3x

OpenStudy (anonymous):

3(x-1)=2x-3+3x 3x -3 = 2x -3 +3x 3x -3 = 5x -3 0 = 2x x =0

OpenStudy (aroub):

3x-3=2x-3+3x 3x-3=5x-3 2x=0 x=0